40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 10 11:55:22 2021, gautam, Update, SUS, PRM suspension suspect PRM_sensorSpectra.pdfPRM_pkFitNaive.pdfPRM_diagComp.pdf
    Reply  Wed Mar 10 17:35:47 2021, gautam, Update, SUS, Spooky action at a distance PRMtrip.pngPRMtrip_zoom.png
Message ID: 15898     Entry time: Wed Mar 10 17:35:47 2021     In reply to: 15894
Author: gautam 
Type: Update 
Category: SUS 
Subject: Spooky action at a distance 

As I am sitting in the control room, the PRM suspension watchdog tripped again. This time, there is clearly no seismic activity. Yet, the BS suspension also shows a slight disturbance at the same time as the PRM. ITMY shows no perturbation though. My best hypothesis here is that the problem is electrical. In Attachment #1, you can see that all of the Sensors go to -6000 cts (whut?) for ~30 seconds. Zooming in to that segment in Attachment #2, it would appear that the light detected by the LED changed dramatically (went dark?) on all 5 coils. The 4 face coils have the same time constant but the side has a different one, but in any case, this level of light change in half a second is clearly not physical. Then the watchdog trips because this huge apparent motion elicits a kick from the damping loops.

The plots I attach are for the DQed sensor channels, so there is some digital filtering involved. But I confirmed that the signal doesn't go negative if I disable the input to the filter module. So it would seem that the voltage input to the ADC really chanegd polarity, seems unphysical. Could be Satellite Box or whitening electronics I suppose - I think we can exclude bad cabling, as that would just lead to the signals going to 0, whereas it would appear here that they did really change sign (confirmed by looking at the ULPDmon channel, which is digitized by Acromag, which reports -10 V at the time of glitch). But why should the BS care about the PRM electronics going wonky?

In addition to an exorcist, we need functioning electronics!


This optic has been hampering my locking attempts all evening. I switched the PRM and SRM satellite boxes, but then I remembered PRM has the Al foil "hats" to attenuate scattered light. of course the Al foil is conducting and can short the OSEM leads. I put some kapton pieces in between OSEM and foil to try and mitigate this issue but I suppose over time it could have slipped, and is making some intermittent contact, shorting PD anode and cathode (that would explain the PD reporting -10 V instead of some physical value).

If this is the problem we would need a vent to address it. In the daytime I'll measure L and R of the coils to see if the actuator imbalance I reported is also due to the same problem...

Attachment 1: PRMtrip.png  360 kB  | Hide | Hide all
PRMtrip.png
Attachment 2: PRMtrip_zoom.png  259 kB  | Hide | Hide all
PRMtrip_zoom.png
ELOG V3.1.3-